OpenStudy (anonymous):

A kilobyte is 210. A megabyte is 220 bytes. How many kilobytes are in a megabyte? Use the exponent rule for division to find your answer and give your answer in exponential form.

OpenStudy (ajprincess):

\(2^{10}~bytes=1~kilobyte\) \(2^{20}~bytes=1~megabyte\) Exponent rule of division \[\large\frac{ a^b }{ a^c }=a^{(b-c)}\] here a=2 b=20 c=10 Does that help? @emily2013

OpenStudy (anonymous):

yes thank you very mch

OpenStudy (ajprincess):

Welcome:)